An individual earns wage income (W) at the rate of Rs. 30,000 per month. He also owns physical assets 

that were worth Rs. 600,000 at the beginning of the fiscal year but now have a market value of 690,000. 

The individual’s income tax liability is determined under a proportional tax system (with ATR = MTR = 

7.5% ) on the basis of his net worth at the end of the fiscal year. He is also subject to paying wealth tax @ 

2% per annum and a capital gains tax @ 1.5% per annum. Calculate total tax liability of the individual 

when the tax system allows for 

(a) zero inflation, and 

(b) 12% inflation.

On the basis of your calculations, explain what is the correct way of taxing capital gains in the presence of 

inflation?


1
Expert's answer
2021-02-16T05:49:22-0500

(a)zero inflation

=Net income*proportion tax

Net income="30000*12=Rs 360000"

Tax of 7.5%

Income tax liability will be

"=360000*7.5\/100"

"=Rs 27000"

Capital gain tax

Capital gain"=690000-600000=Rs 90000"

Tax of 1.5%

Capital gain tax will be

"=90000*1.5\/100"

"=Rs 1350"

Total wealth tax

Total wealth=net income+net assets

"=360000+690000=Rs 1050000"

Tax 2%

"=1050000*2\/100"

"=Rs 21000"

Total tax liability=income tax + capital gain tax + wealth tax"=27000+1350+21000"

"=Rs 49350"

(b) 12% inflation

Income tax liability

Inflation "=360000*12\/100= Rs 43200"

Taxable income=net income - inflation

"=360000-43200=Rs 316800"

Tax of 7.5%

"=316800*7.5\/100"

"=Rs 23760"

Capital gain tax

Inflation"=90000*12\/100=Rs10800"

Taxable gains=total gains - inflation

"=90000-10800=Rs 79200"

Tax of 1.5%

"=79200*1.5\/100"

"=Rs 1188"

Wealth tax

Inflation"=1050000*12\/100=Rs 126000"

Taxable wealth=net wealth- inflation"=1050000-126000=Rs 924000"

Tax of 2%

"=924000*2\/100"

"=Rs 18480"

Total tax liability=income tax+capital gain tax+wealth tax

"=23760+1188+18480"

"=Rs 43428"